A $12,000 loan is to be amortized for 10 years with quarterly payments of $383.06. If the interest rate is 5%, compounded quarterly, what is the unpaid balance immediately after the sixth payment? (Round your answer to the nearest cent.)
Answer #1
Unpaid balance immediately after the sixth payment = Present value of remaining quarterly payments
= quarterly payment * Present value of annuity of 1
= $       383.06 * 27.56046
= $ 10,557.31
Working:
Present value of annuity of 1 = (1-(1+i)^-n)/i Where,
= (1-(1+0.0125)^-34)/0.0125 i = 5%/4 = 0.0125
= 27.5604564 n = (10*4)-6 = 34